OpenStudy (anonymous):

Use less than, equal to, or greater than to complete the statement. The sum of the measures of the exterior angles of a regular 5-gon, one at each vertex, is __________ the sum of the measures of the exterior angles of a regular 9-gon, one at each vertex.

OpenStudy (anonymous):

The sum of the exterior angles of any polygon is ALWAYS 360 degrees One exterior angle = the sum of the exterior angles/the number of sides; One exterior angle = 360/n; where n = number of sides So you just do the math... The measure of a single exterior angle in the pentagon: 360/5 = 72 degrees The measure of a single exterior angle in the nonagon: 360/9 = 40 degrees 72 is greater than 40, so the answer is "Greater than"